# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v2

@@ 2.5 Generating products for chemical reaction
Taking the product prediction in substitution reaction of polycyclic aromatic hydrocarbons (PAHs) based on Raman spectroscopy as an example (see Methods for details of dataset), there may have three situations.
@@
(1) Spectrum-structure retrieval is for the well-known reactions, i.e., predicting the specific substitution site with the known type of substituent. Due to the limited substitution sites of PAHs, it is possible to retrieve by traversing all possible substitution structures, thereby outputting the structure with the highest spectrum-structure similarity. As shown in Table 1, the Recall@1 of Vib2Mol for benzene, naphthalene, and anthracene reached 99.57%, 99.75%, and 99.14%, respectively, indicating Vib2Mol can nearly perfectly perform spectrum-structure retrieval within the limited search space. Obviously, as prior knowledge decreases, the potential research space significantly increases, making it difficult to traverse all possible structures, and the generation is highly demanded.
@@
(2) Conditional generation is for the partial known reactions, i.e., predicting the type of substituent with the known substitution. The average Recall@1 of Vib2Mol is 98.95% in predicting one unknown substituent, and remains at 96.64% for the case of two unknown substituents. Note that the accuracies here are somewhat inflated. Before performing conditional generation, it is necessary to design certain “blanks” for the model to “fill in”, but Vib2Mol directly replaces the characters at the corresponding positions with “<mask>”. This approach may leak the number of characters to be filled. Although we tried to change the number of characters corresponding to “<mask>”, it is hardly to exhaust all possibilities. Obviously, this shortcoming should to be addressed in the future.
@@
(3) De novo generation is for the new reactions, i.e., predicting a completely unknown molecular structure, including the type of substituents and all substitution sites, simultaneously. Due to the simplicity of the structure, the Recall@1 of benzene (95.32%) is significantly better than that of naphthalene (82.44%) and anthracene (84.91%). The average Recall@1 of the three situations can reach 86.63%. Such a high Recall@1 is mainly because of the limited search space of this case, let alone the real condition predicting by a mixed Raman spectrum.
@@
Therefore, by extracting approximately 15,000 chemical reactions listed in the second World AI4S Prize-Material Science Track[21], we calculated the Raman spectra of reactants and products, and mixed them according to the labeled yields (ignoring the differences in Raman scattering cross-sections of different molecules) to simulate the mixed spectra measured in real condition (see Methods for details).
@@
When the expected product is in the database (Figure S2A), Vib2Mol trained on unmixed spectra (Vib2Mol-unmix) achieved a Recall@1 of 84.87% for spectrum-to-structure retrieval on the unmixed spectrum test set. However, performance dropped significantly to 40.47% on the mixed spectra test set, which is closer to the real-world condition. In contrast, Vib2Mol trained on mixed spectra based on yields (Vib2Mol-mix) achieved a Recall@1 of 82.70%. When the expected product is out of the database (Figure S2B), Vib2Mol-unmix achieved a Recall@1 of only 10.17% for de novo generation on the mixed spectrum test set, while Vib2Mol-mix achieved a Recall@1 of 24.40%. These results clearly demonstrate that introducing yield information significantly enhances Vib2Mol’s ability to annotate mixed Raman spectra.
Considering the uncertainty of yields in real experiments, we replaced the fixed yield with a random value within a certain range and further investigated the effect of the range of yield values on performance (Figure S2). The comparable Recall@1 to that with labeled yields implies the retrieval performance is not sensitive to the sampling range of yields. Differently, thanks to the increased diversity of training data with the expanded sampling range, this strategy improves the performance of generation.
It is worth noting that the model trained by unmixed spectra has a Recall@1 of 25.55% for de novo generation on the unmixed spectrum test set, while the model trained with random yields of 10-100% achieved a Recall@1 of 27.65% on the mixed spectrum test set. Such comparison indicates that mixing spectra can slightly enhance generation capabilities of Vib2Mol. The reason for this phenomenon is currently unclear. One possibility is the strengthened characteristic peaks by mixing Raman spectra of reactants (Figure S3), thereby assisting in the generation of product structures.
@@ 2.6 Peptide sequencing and PTMs identification
The Vib2Mol pre-trained on VB-mols was fine-tuned by peptides represented by SMILES. As shown in Table 1, the Recall@1 of the model (Vib2Mol-SMILES) for spectrum-structure retrieval is 63.11%, when the to-be-determined peptide is in the database. Otherwise, the Recall@1 for de novo generation drops to 27.55%. The low Recall@1 ignited us to change peptide representation from SMILES to residue sequences, considering the relatively patterned residue structure. The obtained Vib2Mol-sequence model significantly improved Recall@1 for retrieval and generation up to 67.58% and 39.92%, respectively. This improvement is mainly because of the drastically reduced token length by residue sequences, thereby reducing the complexity and improving the accuracy of sequence generation. This is the reason why current models[3, 12, 26] are mainly based on residue sequences. It is not surprising that the Vib2Mol sequence performed best in elucidating dipeptides (94.74% and 89.47% for retrieval and generation, respectively). Since the search space increases exponentially with the number of residues, the performance of Vib2Mol-sequence gradually decreases with the increasing length of peptides. Nevertheless, even for tetrapeptides, the Recall@1 for retrieval and generation can still reach 65.48% and 36.62%, respectively.
As for the identification of PTMs sites, we constructed the VB-peptide-mod dataset, which includes the most representative phosphorylation and sulfation (see Methods for details), and fine-tuned Vib2Mol-sequence by it. Depending on the level of prior knowledge in practical application, there are three cases (Table 2).
(1) Determining modification type at a specific residue site. Vib2Mol can achieve a high accuracy (75.03%) for the three categories (sulfation, phosphorylation, and unmodified) by conditional generation. (2) Retrieval of peptides within the database. Vib2Mol can reach a Recall@1 of 65.03% by calculating spectrum-structure similarities. (3) De novo sequencing for peptides outside the database. A Recall@1 of 27.99% can be achieved by generation module. As a proof-of-concept, Vib2Mol demonstrates the feasibility of using theoretical Raman spectra in de novo sequencing of peptides and identifying PTMs sites. This advancement holds significant promise for applications in biomedicine, immunology, and drug development. We anticipate a synergistic integration with experimental data will further enhance its utility and uncover new insights.
@@
Vib2Mol-SMILES 63.11 88.13 27.55 33.87
Vib2Mol-sequence 67.58 89.89 39.92 51.18
Dipeptide 94.74 100 89.47 94.74
Tripeptide 77.56 94.45 55.24 68.31
Tetrapeptide 65.48 88.95 36.62 52.51
@@
Accuracy Recall@1 Recall@3 Recall@1
Unmodified 79.16 76.02 93.05 27.26
Phosphorylated 74.97 62.31 86.54 28.94
Sulfated 71.94 58.71 84.68 27.86
Averaged 75.03 65.03 87.77 27.99
2.7 Generalization of Vib2Mol for experimental spectra and multimodal spectral analysis
The above two applications demonstrate the power of Vib2Mol in spectrum-to-structure tasks based on theoretical Raman spectroscopy from chemistry to biology. As it is well-known that there is a huge gap between theoretical and experimental spectra, due to the difficulty in theoretically simulating the interactions among different vibrational modes. Therefore, we explored the feasibility of Vib2Mol in experimental infrared spectroscopy (NIST, containing about  12,000 experimental infrared spectrum-molecular structure pairs), considering the current lack of a large-scale experimental Raman spectroscopy database for molecular elucidation.
As expected, due to the significant differences in spectral features between theoretical and experimental spectra (Figure S4), the Vib2Mol model trained on theoretical spectra performed poorly on experimental database. As shown in Table 3, the Recall@1 for spectrum-structure retrieval and de novo generation were almost 0%, with only conditional generation achieving an accuracy of about 25%. The certain level of accuracy of conditional generation is possibly due to the contextual information of the unmasked structure. When Vib2Mol was trained on experimental infrared spectra, the Recall@1 for spectrum-structure retrieval increased to 19.73%, the accuracy for conditional generation increased to 48.89%, and the Recall@1 for de novo generation increased to 4.41%. Considering the distance is far away from the requirements of practical applications, we pre-trained Vib2Mol on theoretical database, and then employed experimental infrared spectra to fine-tune, resulting in a significant improved performance. The Recall@1 for spectrum-structure retrieval increased to 51.49%, the accuracy for conditional generation increased to 71.55%, and the Recall@1 for de novo generation increased to 11.79%. This improvement is likely due to a certain degree of commonality between theoretical and experimental spectra. Therefore, the pre-training by theoretical spectra provides some assistance in interpreting experimental spectra. No wondering, the performance could be enhanced by introducing other molecular information, such as the chemical formula[9, 14, 27], which is out of the scope of current discussion.
On the basis of the theoretical infrared and Raman databases, we further examined the impact of integrating multi-modal spectral information on Vib2Mol performance. As shown in Table 4, on the spectrum-structure retrieval of VB-mols, there is no much difference in performance between models trained on single-spectral and dual-spectral information. Among them, Raman-only performs the best, with a Recall@1 of 78.92%, while the performance of utilizing Raman+IR is slightly worse (78.83%). One possible reason is that retrieval aims to focus on clear, unique, and highly discriminative features, in order to accurately find the entry in a large number of known molecular structures that best matches the given spectrum. However, Infrared and Raman spectra, observing molecular vibrations from different perspectives, are not completely aligned and unbiased. Therefore, there may be some inconsistency in certain vibrational features, which may have a negative impact on the integration[28]. To reduce this negative impact, one could consider introducing other spectral information from different aspects, such as NMR and MS, to provide more perspectives for a more comprehensive observation of molecular structure features.
On conditional generation and de novo generation, the synergistic effect of the two spectroscopies is very convincing. Taking the de novo generation on VB-mols as an example, the Recall@1 for Raman-only was 59.51%, and the introduction of infrared spectra increased it by 5.5% to 65.01%. This improvement likely stems from the fact that generation tasks prioritize feature comprehensiveness and correlation over highly discriminative features. The diverse molecular structure information provided by different spectra can offer the model richer clues and a more integrate framework for constructing molecular structures or spectra. Therefore, even if some features are not precise enough, the model can still generate reasonably based on the overall information, resulting in better performance.
Table 3: Performance of models trained with different data set tested on NIST-based experimental infrared spectral dataset.
Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Theo. IR 0.1 25.61 0
Exp. IR 19.73 48.89 4.41
Theo. IR + Exp. IR
(finetune) 51.49 71.55 11.79
Table 4: Testing performance of models trained with different datasets.
qm9 zinc15 mols qm9 zinc15 mols qm9 zinc15 mols
Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Theo. IR 80.19 76.02 77.66 97.95 86.23 92.5 57.58 45.16 53.98
Theo. Raman 81.51 76.65 78.92 97.91 87.04 92.88 63.31 49.86 59.41
Theo. Raman + IR 81.03 77.39 78.83 98.35 88.65 93.85 69.38 54.29 65.01
* The “qm9”,“zinc15” and “mols” here all refer to specific subsets within the ViBench.
@@ 3 Discussion
In this study, we proposed Vib2Mol, a DL model for vibrational spectroscopy, which can effectively address multiple spectrum-to-structure tasks according to available prior knowledge. On ViBench, Vib2Mol outperformed traditional methods on 9 of 10 test sets. Such outstanding performance stems from the synergistic of retrieval and generation modules which lead to the better establishment of the correlation between spectrum and molecular structure, which was confirmed by effective clustering molecules with similar structures.
Vib2Mol has shown substantial potential in chemical and biological applications, where we have further tackled several unexplored challenges with in-silico data. On the one hand, chemical reactions inevitably lead to a mixture of reactants and products, which thus results in mixed spectra posing a challenging issue for spectral annotation. We showcased the capability of Vib2Mol to interpret mixed-spectra, which achieved the recall@1 of 82.70% and 24.40% for retrieval and de novo generation on chemical reaction dataset with real yields, respectively. On the other hand, Vib2Mol enables Raman spectroscopy as a unique omics method which not only achieved a Recall@1 of 39.9% for de novo peptide sequencing, but also efficiently predicted PTMs sites of phosphorylated and sulfated modification, where the traditional mass spectrometry falls short. After fine-tuned by experimental spectra, Vib2Mol can be adapted to interpret experimental infrared spectra. Finally it demonstrated potential for integrating multi-modal spectra from multiple techniques as well.
Vib2Mol demonstrates the potential for in situ monitoring of dynamic chemical reactions and life processes on the basis of vibrational spectroscopy. In the future, to better elucidate molecular conformations in dynamic processes, a possible improvement lies in the introduction of stereochemical information. In addition, it is also of great interest to design more flexible generative modules to equip the models with bidirectional spectrum-to-structure and structure-to-spectrum predictions.
@@ 4.1 Reference data
We have established a vibrational spectrum-to-structure benchmark (ViBench, VB), which consists of eight parts: VB-qm9, VB-zinc15, VB-mols, VB-geometry, VB-PAHs, VB-RXN, VB-peptide, and VB-peptide-mod. Details are listed in supplementary information.
Density functional theory (DFT) was employed to perform conformational optimization of these molecules and calculated the corresponding infrared and Raman spectra. Unless otherwise specified, all quantum chemical calculations were carried out using the Gaussian 09 program[g09]. The geometries were optimized using the B3LYP-D3BJ functional with a 6-311+G** basis set. Frequency calculations were obtained at the same level at the optimized geometry.
Furthermore, to test the generalization of Vib2Mol on experimental spectra, we collected experimentally measured infrared spectra of 12,937 small molecules from the public NIST dataset[linstrom1997nist]. For details on the specific division of the above datasets for training, validation, and testing, please refer to Table S4. To facilitate subsequent calculations, the spectral dimensions were unified to 1024, and molecular structures were all represented using SMILES.
@@ 4.2 Related works and baseline models
To fairly compare the performance among Vib2Mol and current methods, we surveyed spectrum-to-structure models based on vibrational spectroscopy in the community. For spectral-structure retrieval, CL is currently the most popular framework. DeepSearch[3], CReSS[5] and SMEN[29] all employ CL to bring the spectra and corresponding structures of the same molecule closer together, achieving great spectral-structure retrieval performance for mass spectrometry, NMR, and IR, respectively. We have borrowed the architecture of the above methods thus building a similar baseline model (vanilla-CL) applicable to Raman spectra. It is worth noting that although SMEN is very similar to Vib2Mol, the two are not comparable because SMEN uses atomic coordinates while Vib2Mol uses SMILES to represent the molecular structure. But in Section 2.3, we have confirmed that the performance of Vib2Mol is better than CL+LM, which is similar to SMEN, in both retrieval and generation.
For the conditional generation, MLM is currently the most popular approach. CO-BERT realized bidirectional prediction between molecular structures and vibrational spectra by MLM[30]. However, CO-BERT focus on predicting atomic coordinates by vibrational spectra and contextual structural information, which cannot be compared with Vib2Mol. Therefore, we followed its architecture and built a similar baseline model (vanilla-MLM).
For the de novo generation, Alberts et al.[14] and Wu et al.[9] both adopted an encoder-decoder architecture and introduced the constraint of molecular formula as well. However, in many cases, information such as the molecular formula is not readily available[7]. Therefore, we followed the above two models and built a formula-free baseline model (vanilla-LM).

### v3

@@ 2.5 Generating products for chemical reaction
Taking the product prediction in substitution reaction of polycyclic aromatic hydrocarbons (PAHs) based on Raman spectroscopy as an example, there may have three situations (Figure 5A).
@@
(1) Spectrum-structure retrieval is for the well-known reactions, i.e., predicting the specific substitution site with the known type of substituent. Due to the limited substitution sites of PAHs, it is possible to retrieve by traversing all possible substitution structures, thereby outputting the structure with the highest spectrum-structure similarity. As shown in Figure 5B, the Recall@1 of Vib2Mol for benzene, naphthalene, and anthracene reached 100.00%, 98.25%, and 99.57%, respectively, indicating Vib2Mol can nearly perfectly perform spectrum-structure retrieval within the limited search space. Obviously, as prior knowledge decreases, the potential research space significantly increases, making it difficult to traverse all possible structures, and the generation is highly demanded.
@@
(2) Conditional generation is for the partial known reactions, i.e., predicting the type of substituent with the known substitution. The weighted average Recall@1 of Vib2Mol is 98.95% in predicting unknown substituent. Note that the accuracies here are somewhat inflated. Before performing conditional generation, it is necessary to design certain "blanks" for the model to "fill in", but Vib2Mol directly replaces the characters at the corresponding positions with "<mask>". This approach may leak the number of characters to be filled. Although we tried to change the number of characters corresponding to "<mask>", it is hardly to exhaust all possibilities. This shortcoming should to be addressed in the future.
@@
(3) De novo generation is for the new reactions, i.e., predicting a completely unknown molecular structure, including the type of substituents and all substitution sites, simultaneously. Due to the simplicity of the structure, the Recall@1 of benzene (98.29%) is significantly better than that of naphthalene (88.80%) and anthracene (88.79%). The weighted average Recall@1 of the three situations can reach 91.39%. The exceptional metrics may primarily be attributed to Vib2Mol’s robust alignment of spectrum-structure, particularly in the constrained generation space of PAHs.
@@
To better reflect real-world conditions, we evaluated Vib2Mol in interpreting mixture spectra of products and reactants. By extracting approximately 15,000 chemical reactions listed in the second World AI4S Prize-Material Science Track[22], we calculated the Raman spectra of reactants and products, and mixed them according to the yields (ignoring the differences in Raman scattering cross-sections of different molecules) to simulate the mixed spectra measured in real condition (see Supplementary Information for details). Spectrum-structure retrieval and de novo generation were used to simulate the scenarios where the expected product is either present in or absent from the database, respectively.
@@
Figure 5B illustrates that Vib2Mol achieved a Recall@1 of 98.11% and 55.84% on the unmixed spectrum test set for retrieval and generation tasks, respectively. However, the performance dropped considerably to 74.29% and 36.81% when tested on mixed spectra, highlighting that training on unmixed dataset is not suitable for mixture analysis for chemical reactions. Therefore, we trained a Vib2Mol-RXN for mixture spectra of products and reactants. Vib2Mol-RXN achieved not only a comparable performance on the unmixed spectral set, but also a significant improvement on the mixed spectrum test set (97.51% for retrieval and 56.64% for generation). These results clearly demonstrate that introducing yield information significantly enhances Vib2Mol’s ability to annotate mixed Raman spectra.
Figure 5: Workflow and performance of Vib2Mol in product elucidation and mixed-spectrum analysis. (A) Three scenarios for predicting products. (B) Benchmarking on PAH substitution reactions. (C) Retrieval and de novo generation results on unmixed and mixed spectra of general chemical reactions.
@@ 2.6 Peptide sequencing and PTMs identification
The Vib2Mol pre-trained on VB-Mols was fine-tuned by peptides represented by SMILES. As shown in Table 1, the Recall@1 of the model (Vib2Mol-SMILES) for spectrum-structure retrieval is 62.97%, when the to-be-determined peptide is in the database. Otherwise, the Recall@1 for de novo generation drops to 50.07%. The low Recall@1 ignited us to change peptide representation from SMILES to residue sequences, considering the relatively patterned residue structure. The obtained Vib2Mol-sequence model significantly improved Recall@1 for retrieval and generation up to 70.33% and 51.49%, respectively. This improvement is mainly because of the drastically reduced token length by residue sequences, thereby reducing the complexity and improving the accuracy of sequence generation. This is the reason why current models[3, 12, 27] are mainly based on residue sequences. It is not surprising that the Vib2Mol sequence performed best in elucidating dipeptides (97.37% and 86.84% for retrieval and generation, respectively). Since the search space increases exponentially with the number of residues, the performance of Vib2Mol-sequence gradually decreases with the increasing length of peptides. Nevertheless, even for tetrapeptides, the Recall@1 for retrieval and generation can still reach 68.03% and 48.18%, respectively.
As for the identification of PTMs sites, we constructed the peptide-mod dataset, which includes the most representative phosphorylation and sulfation (see Methods for details), and fine-tuned Vib2Mol-sequence by it. Depending on the level of prior knowledge in practical application, there are three cases (Table 2).
(1) Determining modification type at a specific residue site. Vib2Mol can achieve a high accuracy (74.43%) for the three categories (sulfation, phosphorylation, and unmodified) by conditional generation. (2) Retrieval of peptides within the database. Vib2Mol can reach a Recall@1 of 65.23% by calculating spectrum-structure similarities. (3) De novo sequencing for peptides outside the database. A Recall@1 of 33.33% can be achieved by generation module. As a proof-of-concept, Vib2Mol demonstrates the feasibility of using theoretical Raman spectra in de novo sequencing of peptides and identifying PTMs sites. This advancement holds significant promise for applications in biomedicine, immunology, and drug development. We anticipate a synergistic integration with experimental data will further enhance its utility and uncover new insights.
@@
Vib2Mol-SMILES 62.97 88.79 50.07 60.66
Vib2Mol-sequence 70.33 90.71 51.49 68.25
Dipeptide 97.37 100.00 86.84 94.74
Tripeptide 81.38 95.93 67.57 82.61
Tetrapeptide 68.03 89.66 48.18 65.34
@@
Accuracy Recall@1 Recall@3 Recall@1 Recall@3
Unmodified 73.79 74.57 91.22 32.90 45.74
Phosphorylated 76.04 62.12 84.79 35.67 47.24
Sulfated 73.73 59.70 82.69 31.94 46.37
Averaged 74.43 65.23 85.90 33.33 46.44
@@ 3 Discussion
In this study, we proposed Vib2Mol, a DL model for vibrational spectroscopy, which can effectively address multiple spectrum-to-structure tasks according to available prior knowledge. It not only achieves state-of-the-art performance in analyzing theoretical Infrared and Raman spectra, but also outperform previous models at experimental data. Such outstanding performance stems from the synergistic of retrieval and generation modules which lead to the better establishment of the correlation between spectrum and molecular structure.
Vib2Mol has shown substantial potential in chemical and biological applications, where we have further tackled several unexplored challenges with in-silico data. On the one hand, chemical reactions inevitably lead to a mixture of reactants and products, which thus results in mixed spectra posing a challenging issue for spectral annotation. We showcased the capability of Vib2Mol to interpret mixed-spectra, which achieved the recall@1 of 97.51% and 56.64% for retrieval and de novo generation on chemical reaction dataset with real yields, respectively. On the other hand, Vib2Mol not only achieved a Recall@1 of 39.9% for de novo peptide sequencing, but also efficiently predicted PTMs sites of phosphorylated and sulfated modification, where the traditional mass spectrometry falls short, which enables Raman spectroscopy as a unique and promising omics method.
Vib2Mol also demonstrates the potential for in situ monitoring of dynamic chemical reactions and life processes on the basis of vibrational spectroscopy. In the future, to better elucidate molecular conformations in dynamic processes, a possible improvement lies in the introduction of stereochemical information. In addition, it is also of great interest to design more flexible generative modules to equip the models with bidirectional spectrum-to-structure and structure-to-spectrum predictions.
@@ 4.1 Reference data
We have developed a vibrational spectrum-to-structure benchmark (ViBench, VB), which consists of two parts: VB-Mols and VB-geometry (VB-GEOM), which can be employed as a lead board for fairly comparing current advanced deep learning models. Additionally, five more benchmarks are established using publicly available datasets to further enhance the reliability of our evaluation. In total, these benchmarks comprise 662,668 theoretical and 17,774 experimental spectra.
From the perspective of data source, these benchmarks are categorized into to theoretical and experimental benchmarks. Theoretical benchmarks include QM9S, VB-Mols, and VB-GEOM. These datasets provide simulated infrared and Raman spectra along with optimized molecular conformations. Among them, QM9S serves as the most basic and widely used theoretical benchmark. It consists of small organic molecules (with fewer than 9 heavy atoms) and their corresponding IR and Raman spectra. The molecules in this dataset are composed exclusively of C, H, O, N, and F elements. Extensionally, ViBench incorporates additional Cl, Br, P, and Si elements, and increases the maximum number of heavy atoms to 45. Experimental benchmarks comprise SDBS (Spectral Database for Organic Compounds) and NIST-IR (United States National Institute for Science and Technology). It is important to note that only SDBS offers both infrared and Raman spectra simultaneously, whereas NIST-IR only provides infrared spectra. Furthermore, molecular conformation information is absent from these experimental datasets.
From the perspective of validation methodology, VB-GEOM is a unique benchmark. It comprises 6,835 molecules, each containing two distinct stable conformers, with each conformer having its corresponding IR and Raman spectra. Thus, each data entry in VB-GEOM is effectively a tuple including 7 elements: SMILES, Conformer#1, IR of Conformer#1, Raman of Conformer#1, Conformer#2, IR of Conformer#2, Raman of Conformer#2. Taking Raman spectroscopy as an example, we treat the spectrum of Conformer#1 as the query and Conformer#2’s as the reference. For retrieval tasks, the model must accurately identify the reference corresponding to the input query, excluding interference from other similar spectra. For generative tasks, the model needs to generate a complete molecular structure based on the input query. The data source, GEOM, is distinct from the QM9 and ZINC15 datasets, and all evaluations are performed under zero-shot conditions.
Moreover, to demonstrate the promising potential of our model for chemical reaction product prediction and peptide sequencing, VB-PAHs, VB-RXN, VB-peptide, and VB-peptide-mod are developed. Details are listed in supplementary information.
Density functional theory (DFT) was employed to perform conformational optimization of molecules and calculated the corresponding infrared and Raman spectra in VB. Unless otherwise specified, all quantum chemical calculations were carried out using the Gaussian 09 program[28]. The geometries were optimized using the B3LYP-D3BJ functional with a 6-311+G** basis set. Frequency calculations were obtained at the same level at the optimized geometry.
For details on the specific division of the above datasets for training, validation, and testing, please refer to Table S6. To facilitate subsequent calculations, the spectral dimensions were unified to 1024, and molecular structures were all represented using SMILES.
@@ 4.2 Related works and baseline models
To fairly compare the performance among Vib2Mol and current methods, we surveyed spectrum-to-structure models based on vibrational spectroscopy in the community. For spectral-structure retrieval, CL is currently the most popular framework. DeepSearch[3], CReSS[5], SMEN[29] and vibraCLIP[19] all employ CL to bring the spectra and corresponding structures of the same molecule closer together, achieving great spectral-structure retrieval performance for mass spectrometry, NMR, IR, and Raman respectively. We fully followed the training and inference configurations of SMEN and vibraCLIP to test their performance on the aforementioned benchmarks. It is worth noting that SMEN, vibraCLIP, and Vib2Mol represent molecules with coordinates (conformation), molecular graphs, and SMILES, respectively. Conformation offers a significant advantage in theoretical spectrum-structure retrieval because a precisely corresponding conformation provides a better molecular representation. However, accurate conformations are unavailable for experimental benchmarks, making only conformation-independent molecular representations easily transferable.
For conditional generation, MLM is currently the most popular approach. CO-BERT realized bidirectional prediction between molecular structures and vibrational spectra by MLM[30]. However, CO-BERT focus on predicting atomic coordinates by vibrational spectra and contextual structural information. Therefore, we built a similar variant based on BERT and compared it with Vib2Mol in Supplementary Information.
For de novo generation, both IR2Mol[14] and the Patch-based Self-Attention (PBSA) model[9] employ an encoder-decoder architecture and integrate molecular formula constraints. We followed the established training and inference protocols of IR2Mol and PBSA to assess their performance across the benchmarks. It is important to note that both of these baselines employ sophisticated data augmentation strategies, and PBSA additionally uses an ensemble learning approach. While we acknowledge the efficacy of these strategies, for the sake of fairness and efficiency in this study, we exclusively used data augmentation applied only to SMILES strings and did not incorporate any ensemble learning strategies in this paper. Recognizing that molecular formulae are not always easily obtainable[7], we developed two distinct versions of Vib2Mol—one that incorporates chemical formulae and another that operates without them. Both versions exhibited strong performance.
Given that datasets like SDBS and NIST-IR are dynamically updated, all benchmarks presented in this paper were constructed using the latest data available as of July 2025. Moreover, to ensure a fresh evaluation, all benchmarks (including QM9S) underwent a complete re-shuffling, rather than adhering to the dataset divisions employed in prior studies.